Assume that the pipe size from the tank to the centrifugal pump is reduced to 0.5 inch.

How would it affect the suction side pressure and operation of the pump?

Just need help with the theory. I believe I need Bernoulli's equation.
Answer #1

By Bernoulli theorem,

P + 1/2 * rho*v2 + rho*g*z = constant

If we this theorem applying into the centrifugal pump and tank then

If we reduced pipe size to 0.5 inch then by continuity equation velocity of pipe increases and when velocity increasing then by Bernoulli theorem suction pressure reduces. Elevation pressure in both case is same

P1 + 1/2 * rho*v12 = P2 + 1/2 * rho*v22

Diameter reduced upto 0.5 inch

Then v2>v1

Then P2<P1

Suction pressure reduced.